### 21 Apr 2026 v1

# Chat2Workflow: A Benchmark for Generating Executable Visual Workflows with Natural Language

Abstract
At present, executable visual workflows have emerged as a mainstream paradigm in real-world industrial deployments, offering strong reliability and controllability. However, in current practice, such workflows are almost entirely constructed through manual engineering: developers must carefully design workflows, write prompts for each step, and repeatedly revise the logic as requirements evolve—making development costly, time-consuming, and error-prone. To study whether large language models can automate this multi-round interaction process, we introduce Chat2Workflow, a benchmark for generating executable visual workflows directly from natural language, and propose a robust agentic framework to mitigate recurrent execution errors. Chat2Workflow is built from a large collection of real-world business workflows, with each instance designed so that the generated workflow can be transformed and directly deployed to practical workflow platforms such as Dify and Coze. Experimental results show that while state-of-the-art language models can often capture high-level intent, they struggle to generate correct, stable, and executable workflows, especially under complex or changing requirements. Although our agentic framework yields up to 5.34% resolve rate gains, the remaining real-world gap positions Chat2Workflow as a foundation for advancing industrial-grade automation.
Chat2Workflow: A Benchmark for Generating Executable Visual Workflows with Natural Language
Yi Zhong♠♡, Buqiang Xu, Yijun Wang, Zifei Shan, Shuofei Qiao, Guozhou Zheng11footnotemark: 1, Ningyu Zhangthanks: Corresponding Author. Zhejiang University  Tencent {zhongyi0212, zhangningyu}@zju.edu.cn https://github.com/zjunlp/Chat2Workflow
1 Introduction
Large language models now form the core of many real-world applications, increasingly deployed through agent-based systems Zhao et al. (2023); Wang et al. (2024); Xi et al. (2025). Existing systems can be broadly divided into ReAct agents Yao et al. (2022) and agentic workflows Zeng et al. (2023); Fan et al. (2024); Qiao et al. (2025). While ReAct agents are flexible, prior work and industry experience suggest that agentic workflows are better suited for reliable and controllable industrial use Shi et al. (2025). Recent interviews Pan et al. (2025) show that simple methods dominate production systems: more than 70% of real-world agent deployments rely on off-the-shelf language models without any weight tuning, instead using explicit workflows for orchestration. This trend is reflected in widely used platforms such as Dify [^1] https://dify.ai/ and Coze [^2] https://www.coze.com/, where agent behavior is primarily predefined through workflows.
Figure 1: An example task in Chat2Workflow, which features realistic, variable natural-language instruction inputs and produces outputs that can be directly transformed and integrated into real-world workflow platforms ( e.g., Dify and Coze).
Despite their prevalence, agentic workflows are still largely created by manual engineering. Human developers must translate natural language requirements into structured workflows, which is time-consuming and error-prone. This raises a natural question: can agentic workflows be automatically generated from natural language? Achieving this is challenging for two reasons. First, real-world requirements are often complex and implicit, making it difficult to infer correct control flow and tool usage from natural language alone. Second, user requirements frequently change, requiring workflows to be revised or regenerated while preserving correctness and consistency.
To study this problem systematically, we introduce Chat2Workflow, the first benchmark focused on generating executable visual workflows from natural language (Figure 1). Chat2Workflow is a high-quality dataset comprising 273 instances in six domains (Figure 2), to evaluate whether generated workflows meet the expected requirements for execution. Each round is equipped with realistic user requests and target workflow components across diverse domains and complexity levels.
The benchmark evaluates whether models can correctly infer workflow structure, select appropriate tools, and produce executable workflows that align with user intent. Experiments on Chat2Workflow show that current state-of-the-art language models can often capture high-level intent but struggle to generate correct and stable workflows, especially as tasks become more complex or requirements evolve. These results indicate that executable workflow generation remains a key bottleneck for practical system automation. Chat2Workflow provides a concrete benchmark for this challenge and points toward future research on structured reasoning and adaptive workflow synthesis for reliable systems.
2 Constructing Benchmark
2.1 Workflow Orchestration Platforms
Overview Dify and Coze are mainstream platforms for building agentic workflows by connecting functional nodes. These workflows are visualized as Directed Acyclic Graphs on the frontend and stored as structured YAML files on the backend. The platforms execute tasks node-by-node, automatically handling tool invocations, inference, and data flow. Taking Dify as a representative example, it provides essential built-in nodes (e.g., Code, If-Else) alongside hundreds of community-driven extensions (e.g., Google Search), enabling the resolution of highly diverse and complex tasks.
Figure 2: Distribution of task types in Chat2Workflow. The benchmark covers six domains: AIGC, Research, Document, Education, Enterprise, and Developer.
Figure 3: Overview of Chat2Workflow benchmark construction and evaluation framework. Left: We collect workflows from six task domains (Research, Document, Enterprise, Developer, Education, AIGC) and reverse-engineer multi-turn instructions. Center: Users interact with LLMs through dialogue, and the model generates workflows in JSON format with CoT reasoning (node selection, design principles, and structured workflow). Right: The JSON workflow is converted to YAML format, uploaded to the platforms for execution, and evaluated against test cases to compute Pass Rate and Resolve Rate.
2.2 Task Formulation
We formalize the generation task of the workflow as a multi-turn interaction process. Given a task instruction for the current round and historical interaction dialogues, we aim to enable language agents to generate graph-structured workflows represented in JSON, which describes the nodes and connectivity in a structured manner. Formally, given the instruction $q$, historical dialogues $H$ and language agents $\mathcal{M}_{\theta}$, the workflow generation can be represented as:
$\mathcal{G}(\mathcal{V},\mathcal{E})\leftarrow\mathcal{M_{\theta}}(q,H),$ (1)
where $\mathcal{G}$ is a Directed Acyclic Graph with node list $\mathcal{V}=\{v_{1},v_{2},...,v_{|\mathcal{V}|}\}$ comprising nodes or tools with predefined parameters populated and edges $\mathcal{E}=\{(v_{i},p,v_{j})\},1\leq i\neq j\leq n,p\in\mathbb{Z}_{\geq 0}$ indicating that the source node $v_{i}$ is connected to the target node $v_{j}$ through the output branch port $p$.
Directly generating executable workflows is challenging for LLMs. To address this, we use a Chain-of-Thought (CoT) approach that outputs a simplified JSON. The model generates selected nodes $\hat{\mathcal{V}}$, planning rationale $p$, and the JSON representation $r$, wrapped in <node_selection>, <design_principle> and <workflow> tags. This JSON is then converted into an executable YAML file. The process is formalized as:
$\mathcal{G}(\mathcal{V},\mathcal{E})\leftarrow r\leftarrow\{\hat{\mathcal{V}},p,r\}\leftarrow\mathcal{M_{\theta}}(q,H),$ (2)
where $\mathcal{V}$ and $\mathcal{E}$ correspond to the nodes_info and edges fields in JSON respectively.
2.3 Dataset Construction
Figure 3 outlines our construction and evaluation framework. To address the absence of multi-turn workflow datasets, we collected production-level Dify and Coze configurations from official sources and GitHub. Based on these workflows derived from actual production, we reverse-engineered the workflow creation instruction set. To ensure that each round’s instruction is based on the previous ones and may add, modify, or refine workflow behavior, we selectively cluster the workflows in the same task context, then rewrite each of them as a round and the whole as a task. To facilitate subsequent evaluation, the written instructions clearly define the input and output variables and give explicit hints about the use of some key nodes. Ultimately, we obtained 27 tasks as the dataset, with each containing 2 to 4 rounds of instructions. 27 tasks can be further classified into the following six types according to the scenarios: Research, Document, Enterprise, Developer, Education and AIGC. More details are in Appendix A.
Although the same task objective can be achieved through different workflows, some specific nodes will inevitably be used, which will be recorded in advance for the corresponding rounds. After the collection and arrangement of the task instructions, we equipped each instruction with three test cases to actually carry out end-to-end testing. For task rounds with relatively clear outputs, the test cases will provide reference answers. For other general cases, only task input is provided. Test cases are primarily derived from existing public datasets, followed by scattered materials from the Internet, and finally synthesized data of language models. The design of test cases takes into account the task requirements of historical rounds, striving to comprehensively cover various situations of multi-branch workflows.
2.4 Data Selection and Processing
We selected the 20 most frequently used nodes to fully cover our dataset. To reduce generation complexity, we simplified the nodes’ I/O variable interfaces: the language agent only needs to provide key-value pairs for essential primary variables, while secondary variables default. In order to endow the language agent with the usage of each node and corresponding I/O interfaces, we craft node knowledge base for 20 node types, serving as a crucial part of the system prompt used for understanding the node ecosystem. On the basis of the above principles, we collect the official node documents, select the specified contents as needed for necessary rewriting, and then organize them into a predefined structured form.
2.5 Quality Control and Dataset Statistics
For the constructed data, we conduct a manual verification process for every individual input–output pair, ensuring the correctness, completeness, privacy and overall quality of the dataset. During this process, workflows that failed to meet our quality standards or violated predefined rules and compliance criteria are systematically identified and removed. This included cases with ambiguous instructions, invalid intermediate steps, time-varying sources, or incoherent workflow behaviors. After this rigorous filtering and validation process, we curate Chat2Workflow, a high-quality dataset consisting of 27 tasks with 79 multi-turn instructions in total, each accompanied by three test instances (totalling 237 instances) to assess whether the generated workflows adhere to the specified execution requirements.
2.6 Evaluation Protocol
Unlike general generation tasks such as text and images, the quality of a workflow cannot be judged entirely based on its representation form itself. It is essential to check whether it can be executed normally and what the execution outcome is. Therefore, we adopted a two-stage progressive evaluation method. The first stage aims at the workflow representation form itself, checking whether it is consistent with CoT sequence and whether the generated workflow JSON can be converted into a valid YAML file. The second stage involves verifying whether the workflow can be successfully imported into the platforms, correctly executed, and produce results that satisfy the specified requirements. We introduce two metrics, pass rate and resolve rate, corresponding to the first and second stages respectively.
Data: Model response $\mathcal{R}$, Reference variables $\mathcal{V}_{ref}$, Key nodes $N_{key}$
Notation: Selected nodes $\hat{\mathcal{V}}$, Planning rationale $p$, JSON file $r$, YAML file $F_{yaml}$}, Workflow $\mathcal{G}$
// Step 1: Format Check
$\hat{\mathcal{V}},p,r\leftarrow\textsc{Parse}(\mathcal{R})$;
if $\mathcal{R}$ lacks required tags or JSON $r$ not parsable then
 return FalseFalse;
// Step 2: Conversion & Import
${F_{yaml}}\leftarrow\textsc{JsonToYaml}(r)$;
if conversion fails then
 return FalseFalse;
$\mathcal{G}\leftarrow\textsc{Import}(F_{yaml})$;
// Step 3: Variable Consistency
if $\textsc{ExtractVars}(r)\neq\mathcal{V}_{ref}$ then
 return FalseFalse;
// Step 4: Logical Validity
if $f_{LLM}(\hat{\mathcal{V}},p,r)$ indicates inconsistency then
 return FalseFalse;
[…]
 return FalseFalse;
return TrueTrue;
Algorithm 1 Pass Rate Pipeline
Pass rate measures the proportion of subtasks whose generated JSON representation is consistent with the CoT sequence as well as the predefined reference variables, and can be successfully converted into valid YAML files, following the sequential checks outlined in Algorithm 1.
If the above four steps are successfully completed, the workflow will be marked as passed and imported into the platforms. It can be calculated through the following formula:
$\text{\%Pas.}=\frac{N_{\text{success\_subtask}}}{N_{\text{total\_subtask}}},$ (3)
where %Pas. indicates Pass rate, $N_{\text{success\_subtask}}$ indicates the number of successful subtasks, $N_{\text{total\_subtask}}$ indicates the total number of subtasks.
Data: Workflow $\mathcal{G}$, instructions $\mathcal{I}$,Test case $\mathcal{T}=\{\text{input},\text{ref\_output}\}$
Notation: Execution output $\mathcal{O}$
// Step 1: Execution Check
$\mathcal{O}\leftarrow\textsc{Execute}(\mathcal{G},\mathcal{T}.\text{input})$;
if execution fails with errors or $\mathcal{O}=\emptyset$ then
 return FalseFalse;
// Step 2: Output Validation
$\mathcal{O}_{file},\mathcal{O}_{text}\leftarrow\textsc{SplitOutput}(\mathcal{O})$;
if $\mathcal{O}_{file}\neq\emptyset$ and file types/extensions mismatch then
 return FalseFalse;
if $\mathcal{O}_{text}=\emptyset$ then
 return TrueTrue;
if $f_{LLM}(\mathcal{I},\mathcal{O}_{text},\mathcal{T}.\text{input},\mathcal{T}.\text{ref\_output})$ confirms all requirements met then
 return TrueTrue;
return FalseFalse;
Algorithm 2 Resolve Rate Pipeline
Resolve rate measures the ratio of test cases where generated workflow can produce the results that meet all specified requirements of instructions. Algorithm 2 details the two-step verification process. Similarly, we will conduct the checks in sequence according to the process. It can be calculated through the following formula:
$\text{\%Res.}=\frac{N_{\text{success\_case}}}{N_{\text{total\_case}}},$ (4)
where %Res. refers to Resolve rate, $N_{\text{success\_case}}$ indicates the number of test cases whose execution results are in line with the requirement, $N_{\text{total\_case}}$ indicates the total number of test cases.
We employ DeepSeek-V3 for evaluation. The corresponding prompts related to the pass and resolve stages are detailed in Figures 6 and 7 of the Appendix, respectively. To assess performance, we conducted independent sampling for both stages, drawing 500 samples for the pass stage and 1,282 samples for the resolve stage. The LLM’s assessments demonstrated exceptional alignment with human evaluation, achieving 100% agreement for the pass stage and 98.83% agreement for the resolve stage.
3 Experiments
3.1 Experimental Setup
To thoroughly assess workflow generation capabilities, we validate 15 representative models on the Chat2Workflow benchmark, reporting the average results across three independent runs:
1) Four closed-source LMs: GPT-{5.1, 5.2}, Claude-Sonnet-4.5, and Gemini-3-Pro-Preview.
2) Eleven open-source LMs: Qwen-3-{8B, 14B, 32B}, Qwen3-235B-A22B, Qwen3-Coder-480B-A35B-Instruct, GLM-{4.6,4.7}, DeepSeek-{V3.1,V3.2}, Kimi-K2-{Instruct, Thinking}.
Furthermore, to explore the performance upper bound, we select two representative models (GPT-5.1 and GPT-5.2) for an advanced agentic evaluation, which is detailed in Section 3.5.
The experiments are conducted on the 1.9.2 version of Dify. More details are in Appendix B.
Model Research Document Enterprise Developer Education AIGC Average
%Pas. %Res. %Pas. %Res. %Pas. %Res. %Pas. %Res. %Pas. %Res. %Pas. %Res. %Pas. %Res.
Closed-Sourced
GPT-5.1 48.89 45.19 53.33 42.22 50.00 33.33 20.83 12.50 36.36 27.27 57.41 55.56 47.26 39.38
GPT-5.2 68.89 60.74 66.67 55.55 86.11 58.33 70.83 45.83 75.76 62.63 48.15 45.68 67.51 54.71
Claude-Sonnet-4.5 68.89 62.97 75.56 64.44 61.11 33.33 75.00 33.33 90.91 76.77 62.97 49.38 71.31 54.57
Gemini-3-Pro-Preview 77.78 77.78 82.22 77.04 86.11 75.93 70.83 52.78 90.91 72.73 74.08 66.67 80.17 71.59
Open-Sourced
Qwen-3-8B 40.00 21.48 8.89 4.45 8.33 0.00 12.5 2.78 3.03 0.00 7.41 7.41 13.92 6.89
Qwen-3-14B 26.67 21.48 33.33 17.78 13.89 0.00 20.83 12.50 27.27 5.05 35.18 25.31 27.43 15.19
Qwen-3-32B 31.11 26.67 28.89 14.07 30.56 8.33 50.00 27.78 33.33 28.28 48.15 33.34 36.71 23.35
Qwen-3-235B-A22B 35.55 32.59 51.11 38.52 19.45 2.78 62.50 29.17 39.39 32.32 42.59 27.78 40.93 27.71
Qwen-3-Coder-480B-A35B-Instruct 53.33 39.26 57.78 40.00 25.00 1.85 25.00 0.00 45.45 26.26 37.03 32.71 42.19 26.44
GLM-4.6 62.22 60.00 60.00 45.93 63.89 24.07 45.83 6.94 72.73 29.29 64.82 56.79 62.45 41.49
GLM-4.7 80.00 60.00 62.22 54.81 72.22 47.22 41.67 29.17 78.79 65.66 70.37 65.43 69.20 55.98
Deepseek-V3.1 51.11 51.11 51.11 36.29 61.11 26.85 20.83 4.17 45.45 39.39 62.96 55.56 51.48 39.24
Deepseek-V3.2 51.11 43.70 46.67 36.30 27.78 4.63 16.67 4.17 42.42 29.29 57.41 53.09 43.46 32.49
Kimi-K2-Instruct 51.11 48.89 37.78 33.33 11.11 8.33 29.17 9.72 21.21 15.15 51.85 48.76 36.29 31.08
Kimi-K2-Thinking 55.56 50.37 66.67 52.59 52.78 25.93 62.50 33.33 27.27 19.19 72.22 61.11 57.81 43.46
Table 1: Experimental results across six task domains for 4 closed-source and 11 open-source models. Pass Rate (%Pas.) measures format correctness, while Resolve Rate (%Res.) measures actual problem-solving capability.
Figure 4: Performance degradation across dialogue rounds. We show the Pass Rate and Resolve Rate for all 15 models across the first three dialogue rounds. Most models exhibit a steady decline in both metrics as the number of interaction rounds increases, indicating the challenge of maintaining workflow quality under evolving requirements.
Figure 5: Bad case analysis for the StudyPlanner task. We compare outputs from three representative models: Kimi-K2-Instruct produces an invalid edge connection in the iteration node; GPT-5.2 generates a correct and executable workflow; GLM-4.6 outputs logically inconsistent node declarations. This illustrates common failure modes in workflow generation.
3.2 Main Results
Table 1 presents our experimental results, which we use to analyze the following research questions.
Q1: What is the gap between the workflow pass rate and resolve rate? For each kind of task, we show both the pass rate %Pas. and resolve rate %Res., representing the format legality and practical usability of workflows respectively. It can be observed that the resolve rate of all models is lower than the pass rate, with the largest disparity observed in GLM-4.6, reaching an average difference of 20.96%. Although Kimi-K2-Instruct shows the lowest gap of 5.21%, both of its absolute values are at a low level compared to the highest visible resolve rate of 71.59%. In addition, in Developer scenario, it has also shown a significant decline. In terms of specific task scenarios, the gap of GLM-4.6 in Education task scenarios has reached an astonishing 43.44%. We conclude that a legal workflow representation that meets the format requirements is still far from successful execution. The former merely serves as a fundamental and necessary condition, but not a sufficient one for achieving the latter, which often has the phenomenon of inflated scores.
Q2: To what extent do existing LLM agents approach the capabilities of real workflow design experts? Since the seed dataset is collected from typical workflows commmonly used by people, together with the mechanism for creating executable workflows through interactive dialogues and the requirements for using files as input and output, our benchmark can to a large extent represent the real-world scenarios. Considering whether the workflow can solve the ultimate problem only, Gemini-3-Pro-Preview performs the best on average, with its absolute performance of 71.59%, which is still far from the level of a practical and deployable workflow planner. Even the best open-source model GLM-4.7 can achieve 55.98% in resolve rate, it still struggles significantly in the Developer scenario. Given the limited number of node types considered, real-world workflow creation is inherently more complex; consequently, Gemini-3-Pro-Preview is supposed to show inadequate performance on complex workflows involving a larger set of nodes.
Q3: How can we improve the workflow generation capabilities of LLM agents? 1) First, by examining the performance of models in the Qwen-3 series, we can evaluate the impact of model size on workflow generation. From the perspective of average indicator, the performance improvement has been stable and significant as the number of model parameters has increased from 8B to 32B and then to 235B. We can draw a conclusion that a larger model size is conducive to workflow generation. 2) Second, post-training is a commonly used method to enhance models in a specific domain. In this domain, it can indeed make the generated workflow more in line with format specifications. However, a high pass rate does not necessarily lead to a high resolve rate. For example, in the Enterprise scenario GLM-4.6 has a higher pass rate than GLM-4.7, with 45.83% against 41.67%. But in terms of the resolve rate, GLM-4.7 is actually higher, with 29.17% against 6.94% Therefore, if only post-training is used to enhance compliance with the format, it may not be able to effectively improve the problem-solving ability of the workflow. 3) Last but not least, it is observed that the performance of instruct model is worse than think model. With larger model size, Qwen-3-Coder-480B-A35B-Instruct has a lower resolve rate than Qwen-3-235B-A22B, with 26.44% against 27.71%. Similarly, as both belong to the Kimi-K2 series, Kimi-K2-Thinking consistently outperforms the instruct model in all task domains. So the thinking mechanism can effectively promote the workflow generation capability.
3.3 Analysis of User Requirement Changes from Chat Interactions
Each task goes through 2 to 4 rounds of dialogue, and each round is regarded as a subtask. The workflow generation requirements for each round will be based on historical conversations, with additions, deletions or modifications made to the existing structure, which poses a further challenge to the model’s ability of complex task understanding and long-horizon instruction following.
To prevent skewed representations—as only 4 tasks contain four dialogue rounds—we illustrate the trend of the two evaluation metrics over the first three rounds for each model, where each data point is the average of three experimental runs (Figure 4). It can be intuitively observed from the table that for most models, the quality of the generated workflow shows a stable downward trend as the number of interaction rounds increases, which corroborates the challenges mentioned above. The slight improvement observed for GPT-5.2 model from the second to the third round is due to some tasks with only two interaction rounds, which affect the second round but not the third.
Furthermore, it can be observed that in most models, a common phenomenon occurs that as the rounds progress, the absolute value of the slope corresponding to the broken line gradually decreases. It might be explained as diminishing marginal returns. The workflows that can still be legally parsed and successfully executed in the second round have already undergone one screening, and they possess a certain ability to understand changes in workflow requirements. With more rounds of modifications, the rate of performance degradation would be reduced, reflecting the distinction between qualitative changes and cumulative quantitative effects.
3.4 Case Analysis
We take the StudyPlanner task of Education scenario as an example for analysis. In the first two rounds, the given instructions are as follows.
Round 1: Create a learning path planning workflow. The user will provide a descriptive instruction (variable instruction) as input. The task is to extract four core fields from the input: interested learning field, learning goals, learning preferences, and average learning duration. Subsequently, based on the extracted information, provide the user with a detailed self-study improvement plan, requiring Markdown format output. The workflow needs to output the self-study plan (variable plan).
Round 2: Modify based on the existing foundation. The workflow needs to be able to automatically generate a full set of tutorials. It requires first generating a course syllabus, and then iteratively generating knowledge points by chapter. The content is required to be rigorous, containing rich examples, pros and cons, and precautions. The workflow only needs to output the final tutorial integrated through template conversion (variable tutorial).
In the second round, the responses given by different models are shown in the Figure 5. Specifically, the visual workflow generated by GPT-5.2 within the Dify and Coze platforms are illustrated respectively in Figures 8 and 9 of the Appendix. 1) Kimi-K2-Instruct outputs an incorrect edge connection relationship. The node with id "5" is an "iteration" node, which internally contains a sub-workflow starting from the "iteration-start". However, there is only an inclusion relationship between them, and no edge connection relationship. This point has been emphasized in the node documentation. Therefore, the generated workflow can not really run. 2) GPT-5.2 creates a legal and executable workflow, while successfully resolving the given case. 3) There are logical inconsistencies in the content output by GLM-4.6. The node "code" and "iteration-start" are present in the workflow but not declared in <node_selection>, which implies the model’s grasp of workflow creation logic is flawed and incoherent in this instance, rendering the final workflow practically unusable.
3.5 An Error-Driven Agentic Baseline
Model Pass Rate Resolve Rate
Zero-Shot Agentic Zero-Shot Agentic
GPT-5.1 47.26 64.14 39.38 44.31
GPT-5.2 67.51 78.06 54.71 60.05
Table 2: Performance comparison of GPT-{5.1, 5.2} under Zero-Shot and the proposed Agentic settings.
To address the bottlenecks identified in Section 3.4 and beyond, we propose an error-driven agentic baseline to explore the performance upper bound of LLMs on the benchmark. Implemented via the OpenCode framework, this approach transitions from vanilla zero-shot prompting to a structured SKILL-based paradigm, explicitly incorporating task-critical guidelines, as well as rules for multi-turn interactions and variable references.
To mitigate context decay in subsequent interaction turns, we dynamically extract Variable Summaries from preceding workflow iterations as supplementary context. Crucially, we construct a robust execution loop featuring a 5-attempt retry mechanism with comprehensive structural and semantic verification. Upon failure, the framework triggers targeted Auto-Repair Modules to rectify the four most prevalent errors: code fence malformation, JSON decoding failures, topological sorting violations, and node selection inconsistencies. As shown in Table 2, the agentic framework yields substantial improvements, boosting Resolve Rates for GPT-5.1 and GPT-5.2 by absolute margins of 4.93% and 5.34%, respectively.
4 Related work
4.1 LLM-based Agents
The rise of LLM has greatly promoted the development of AGI, providing a capability foundation for LLM-centered AI agents. Since LLM-based agents can fully leverage the powerful task understanding capabilities of LLM, they are widely used to solve complex real-world problems with the aid of general tools Schick et al. (2023); Patil et al. (2024), collaborative frameworks Hong et al. (2024); Qian et al. (2023), and extensive API repositories Qin et al. (2024); Tang et al. (2023); Zhang et al. (2024a); Yang et al. (2024). Many prompt-based methods Wei et al. (2022); Yao et al. (2022); Zhu et al. (2026) are proven effective for improving performance. However, these agent methods or frameworks mainly focus on end-to-end effects, paying little attention to the requirements and norms of the intermediate processes in solving tasks, such as planning and reasoning paths, which is not conducive to stable and reliable execution, as well as the reproduction of results across various domains Ge et al. (2024); Xie et al. (2024); Liu et al. (2024); Yang et al. (2026); Liang et al. (2026).
4.2 Workflow and Language Agent Planning
Serving as an intermediate state, workflow provides a way to bridge the gap between task goals and the specific executable steps. On the one hand, by breaking down the task into several executable atomic steps and organizing them into execution plans based on formal logical relationships van der Aalst (1997); Dijkman et al. (2008), the workflow achieves complete transparency of the reasoning path, effectively enhancing the interpretability and reliability of the results. On the other hand, the introduction of prior organizational structures makes it possible to construct the logic of complex tasks using established patterns van der Aalst et al. (2003); Commoner et al. (1971) and robotic process automation Ivančić et al. (2019); Hofmann et al. (2020). Early workflow relied on meticulous manual design to avoid the hallucinations. However, hand-crafted workflows are time-consuming and labor-intensive. Therefore, many studies have shifted their focus to using LLM for automated workflow generation Shen et al. (2023a); Zeng et al. (2023), including iterative synthesis Zhang et al. (2024b); Li et al. (2024) and knowledge-augmented planning frameworks Ye et al. (2023); Zhu et al. (2024); Wornow et al. (2024); Huang et al. (2024). More related work can be found in Appendix C.
5 Conclusion
We introduce Chat2Workflow, a benchmark for evaluating LLMs’ ability to generate deployable visual workflows from natural language, and show that even when augmented with an agentic framework, current models remain fragile under structural constraints and requirement changes. We hope Chat2Workflow can provide both a realistic testbed and a concrete target for future research on automated workflow engineering, and will help bridge the gap between language model-driven workflow design and industrial deployment.
Limitations
Despite its contributions, Chat2Workflow has several limitations. First, while the high-quality dataset is manually verified , the current scale may not encompass the near-infinite variety of logic found in complex industrial business processes. Second, we simplified node interfaces to prioritize executability, which may not fully capture the intricate parameter configurations required in some real-world deployments. Finally, the current system incorporates only 20 high-frequency node types. These specific nodes were selected as they effectively cover most standard workflow scenarios; however, a wide range of valuable tools remain to be incorporated.
Appendix A Detailed Dataset Description
The description of the task scenarios is as follows:
Research includes 5 tasks, which focus on systematically obtaining, understanding, and integrating knowledge from unstructured data. They typically take papers, books, or research topics as input, emphasizing in-depth reading of the content, logical organization, and critical thinking skills.
Document includes 5 tasks, whose core value lies in transforming raw inputs such as files, images, and tables into structured results that are computable, reusable, and processable downstream. They serve as an important representative of the “AI preprocessing layer” in data processing.
Enterprise includes 4 tasks, which are targeted at real Enterprise scenarios, with the goal of automating, standardizing, and providing decision support for complex business processes through AI. Their inputs are mostly internal or business-related documents from enterprises.
Developer includes 3 tasks, which serve scenarios like software development, system design, and technical understanding. The core objective is to lower the threshold for technical understanding and development and enhance engineering efficiency.
Education includes 4 tasks, which revolve around the complete life cycle of “teaching and learning”, with the goal of building a closed-loop system from learning planning, teaching content generation, assessment to feedback.
AIGC includes 6 tasks, which are positioned at the core of “generating directly usable content or materials from 0 to 1”, emphasizing creative expression and multimodal output capabilities. The input is often highly abstract or brief instructions, while the output is complete, publishable, and consumable content results.
Appendix B Experimental Settings
Dify platform requires the necessary extensions to be prepared in advance. A set of common models is installed to make the workflows runnable and the results reproducible. Workflows in Dify are natively represented in YAML format. We have developed a code framework for rule-based transformation to convert the JSON format workflow generated by the language agent into valid YAML format. This ensures that workflows can be effectively generated by agents through simplified JSON representation while remaining fully compatible with the Dify execution environment. In addition, the CoT reasoning sequence is uniformly used as the output method. The node types included in the knowledge base are as follows: Start, End, LLM, Question Classifier, Code, Document Extractor, HTTP Request, If-Else, List Operator, Parameter Extractor, Template, Variable Aggregator, Iteration, Iteration-Start, Text to Speech, Text to Image, Mermaid Converter, Markdown Exporter, Google Search and Echarts. Specifically, the Question Classifier and Parameter Extractor node types are strictly designated to the Qwen-3-Max (qwen3-max), whereas the LLM node type is fixed to the Qwen-3-VL-Plus (qwen3-vl-plus). Text to Image node type are fixed to use the Z-image-Turbo (z-image-turbo) and Text to Speech node type are fixed to use the GPT-4o-Mini-TTS (gpt-4o-mini-tts). For Qwen-3-VL-Plus, temperature is 0.7, max_tokens is 32768, other hyperparameters used during decoding are all set to default values. The system prompt and SKILL.md for workflow generation will be integrated into the GitHub repository soon. As for the Agentic Baseline method, the opencode framework we use is version 1.3.17.
Appendix C More Related Work on Automated Workflow Evaluation
It is crucial to evaluate the quality of the workflow created by LLM. Previous studies have tried to automate workflow evaluation in tool learning scenarios Li et al. (2023); Shen et al. (2023b) and fine-grained planning tasks Chen et al. (2024); Ye et al. (2024); Zheng et al. (2024), using metrics to assess format validity and tool consistency. But these works still have some limitations: Firstly, Most of them remain at the stage of abstract expression of the workflow Qiao et al. (2024), only checking key elements without considering actual execution effects Guo et al. (2024), structural dependencies Lal et al. (2024), or platform-specific constraints Shen et al. (2024). They may identify obvious format errors but cannot ensure usability. Secondly, these works focus on single-round evaluation, neglecting scenarios where requirements change over time.
You are a strict Dify Workflow evaluator. Your task is to determine whether a workflow design is valid based only on the information provided. Do not make assumptions, infer missing nodes, or introduce any external knowledge. # Inputs I will provide four sections: 1. node_selection: A list or description of nodes selected for the workflow. It represents the nodes the designer claims will be used. 2. design_principle: Design principles or constraints for the Dify workflow. It is used to judge overall logical consistency. 3. workflow: The JSON string representation of the Dify workflow. You must parse this JSON to extract the nodes that are actually used. 4. gt_nodes: A list of mandatory (ground-truth) nodes. These nodes must be included to be considered valid. # Evaluation Rules All rules must be satisfied for the result to be true. If any rule fails, the result must be false. Rule 1: Ground-Truth Node Coverage Check whether node types in gt_nodes is a subset of node_selection. The node type only needs to appear, and there is no need to consider the frequency of occurrence. If any node type in gt_nodes is missing from node_selection, immediately return false. Do not continue further checks. Rule 2: Consistency and Exact Node Matching You must verify all of the following conditions: a. Logical Consistency: node_selection, design_principle, and workflow must be logically consistent. The workflow structure must not violate the stated design_principle. b. Exact Node Set Matching (Bidirectional Constraint): Extract the actual node set from the workflow JSON. This node set must exactly match the nodes declared in node_selection. The node type only needs to appear, and there is no need to consider the frequency of occurrence. Specifically: Nodes not declared in node_selection must not appear in workflow. Nodes declared in node_selection must appear in workflow. **Note: 1. The type of the "Template" node is "template-transform". 2. There is no need to distinguish between uppercase and lowercase letters.** # Final Decision Logic Return true only if Rule 1 and Rule 2 are fully satisfied. If any single condition fails, return false. There is no partial credit or “mostly correct” outcome. # Output Format Your response must follow this exact format and nothing else: <reason> [Explain the evaluation reasoning, stating which rules were satisfied or violated] </reason> <result> [true or false] </result>
Figure 6: Prompt for evaluating pass rate.
You are an expert evaluator of workflow execution quality. Your task is to judge whether the execution result of a workflow satisfies the requirements of the current round instruction, based on all available information. # Inputs You Will Be Given You will be provided with the following four parts: 1. queries: A list of historical instructions for creating or modifying the workflow. - Each round’s instruction is based on the previous ones and may add, modify, or refine workflow behavior. - The latest instruction represents the requirements for the current evaluation round. - Ignore the input and output requirements for the file in the instruction of query. The file part has been evaluated separately. 2. input: The non-file input variables used in the workflow execution of the current round. - This field may be empty. 3. output: The non-file output variables produced by the workflow execution of the current round. 4. reference_answer: A reference answer for the expected result. - It may contain minor omissions. - Its format may not strictly match the required output format. - However, its content is considered correct. - This field may be empty. # Evaluation Instructions Please evaluate whether the workflow execution meets the requirements of the current round instruction, following these principles: 1. Instruction Alignment - Identify the requirements implied or explicitly stated in the latest instruction within queries. - Consider earlier instructions only insofar as they remain valid and are not overridden. 2. Information Utilization - Use all available information (queries, input, output, reference_answer) for judgment. - Both of input and reference_answer may be empty; this alone does not imply failure. 3. Output Correctness Criteria - If output exists, judge whether its content and format satisfy the current instruction, except for the requirements related to the files. - If reference_answer exists, use it as a semantic reference rather than a strict template. 4. Handling Missing or Empty Fields - If outputs clearly contradict the instruction, the result should be false, except for the requirements related to the files. - If reference_answer is empty, rely on instruction compliance and logical consistency only. 5. Final Judgment Rule - If the workflow execution reasonably fulfills the intent and requirements of the current round instruction (except for the requirements related to the files), return true. - Otherwise, return false. **Note: This rule takes precedence over all the above —— The file variable part in the instruction requirements, whether as input or output, including images, documents, and audio will not be provided in this evaluation. So it cannot be used as the basis for a judgment of ’False’!!! For example, if the instruction or query requires the output of a PDF document but the output does not include the document, this point should be ignored at this time to check whether other aspects meet the requirements. ** # Output Format Your response must follow this exact format and nothing else: <reason> [Explain the evaluation reasoning, stating which rules were satisfied or violated] </reason> <result> [true or false] </result>
Figure 7: Prompt for evaluating resolve rate.
Figure 8: The Dify workflow generated by GPT-5.2 in the second round of the Studyplanner task.
Figure 9: The Coze workflow generated by GPT-5.2 in the second round of the Studyplanner task.